OpenStudy (sugarmochi):

2x + 3y = 7 3x + 12y = 9 What is x and y?

OpenStudy (luigi0210):

You could use a sub

OpenStudy (sugarmochi):

sub?

OpenStudy (luigi0210):

Substitution

OpenStudy (sugarmochi):

dont know how..

OpenStudy (luigi0210):

Second equation, you could solve for x

OpenStudy (anonymous):

Does it matter what method we use? Eliminate/ substitution?

OpenStudy (triciaal):

to solve simultaneous equations there are several methods that can be used; graphing, substitution, elimination and matrices. the bottom line is the solution satisfies each equation. for this problem lets use substitution. we will divide the 2nd eqn by 3 since it can be factored and find the equivalent of x then we will substitute for x in the 1st eqn. this gives an equation with just one variable y. we solve for y then put this value back in the expression to find x. It is a good practice to check the solution.

OpenStudy (triciaal):

3x + 12y = 9 divide by 3 gives x + 4y = 3 means x = 3 - 4y substituting in the 1st eqn 2x + 3y = 7 2(3 - 4y) + 3y = 7 6 - 8y + 3y = 7 y = -1/5 x = 3 - 4(-1/5) = (15 + 4)/5 = 19/5 check the ans in eqn 2 3x + 12 y = 9 left hand side 3*(19/5) + 12(-1/5) = (57-12)/5 = 45/5 = 9 = right hand side and therefore correct solution 19/5, -1/5

OpenStudy (triciaal):

to do the same problem by elimination you would multiply the 1st eqn by 4 and subtract the 2nd eqn. this will eliminate the y. solve the one variable eqn for x and then substitute to find y 2x + 3y = 7 multiply by 4 8x + 12 y = 28 subtract 2nd eqn 3x + 12y = 9 5x = 19 gives x = 19/5 you have x find y 3(19/5) + 12 y = 9 57/5 -9 = -12 y divide by -12 so y = -1/5 different method same result

OpenStudy (triciaal):

to graph: plot some points for each line and find where they intersect to use matrices; will not cover that right now you have enough for now

OpenStudy (triciaal):

do you have any questions?

OpenStudy (sugarmochi):

Wow